### v1

# Multi-group Uncertainty Quantification for Long-form Text Generation

Abstract
While large language models are rapidly moving towards consumer-facing applications, they are often still prone to factual errors and hallucinations. In order to reduce the potential harms that may come from these errors, it is important for users to know to what extent they can trust an LLM when it makes a factual claim. To this end, we study the problem of uncertainty quantification of factual correctness in long-form natural language generation. Given some output from a large language model, we study both uncertainty at the level of individual claims contained within the output (via calibration) and uncertainty across the entire output itself (via conformal prediction). Moreover, we invoke multicalibration and multivalid conformal prediction to ensure that such uncertainty guarantees are valid both marginally and across distinct groups of prompts. Using the task of biography generation, we demonstrate empirically that having access to and making use of additional group attributes for each prompt improves both overall and group-wise performance. As the problems of calibration, conformal prediction, and their multi-group counterparts have not been extensively explored previously in the context of long-form text generation, we consider these empirical results to form a benchmark for this setting.
Multi-group Uncertainty Quantification for Long-form Text Generation
Terrance Liu Carnegie Mellon University Pittsburgh, PA 15213 terrancl@cs.cmu.edu                        Zhiwei Steven Wu Carnegie Mellon University Pittsburgh, PA 15213 zstevenwu@cmu.edu
1 Introduction
In recent years, researchers have developed stronger large language models that perform well on a variety of tasks across different domains (Touvron et al., 2023; Bubeck et al., 2023; Jiang et al., 2023; Anil et al., 2023). In particular, researchers have demonstrated the capabilities of such models by showcasing their ability to generate long-form text. However, as use of LLMs continues to grow, so do concerns over their tendency to hallucinate facts (Huang et al., 2023). As a result, there is a growing need for methods that can reduce hallucinations (Manakul et al., 2023; Zhang et al., 2023), perform abstention (Yang et al., 2023), or provide correctness guarantees (Kumar et al., 2023; Mohri and Hashimoto, 2024; Quach et al., 2023).
In our work, we focus on the latter—broadly speaking, uncertainty quantification of long-form large language model generations, which is a problem that has not previously been exhaustively explored. In contrast to the few existing works on factuality in long-form generations (Quach et al., 2023; Mohri and Hashimoto, 2024), we make the observation that while uncertainty guarantees may be valid under the full data distribution, they may not still be valid within individual subgroups of the distribution. For example, generations describing local politicians may be more prone to error than generations concerning national leaders. To address this unmet need, we introduce methods quantifying uncertainty in long-form text generation that are valid not only across a full distribution of prompts (i.e., marginally) but also across identifiable subgroups of prompts (i.e., conditionally).
Concretely, given a set of claims produced by an LLM in response to some prompt, our goal is to provide a confidence score or uncertainty guarantee about the factual correctness of the output. We explore this problem in two settings. Given a set of claims contained within some long-form prompt response, we (1) ensure factuality at the individual claim level and (2) provide uncertainty guarantees across the whole set of claims. We approach problem (1) via (multi)calibration (Hébert-Johnson et al., 2018), in which one wishes to output a calibrated score for each claim, while for problem (2), we apply (multivalid) conformal prediction (Jung et al., 2022), selecting some subset of claims that with high probability are all factually correct.
Empirically, we evaluate our methods using biography generation as a testbed for multi-group uncertainty quantification of factuality. As the problems of calibration, conformal prediction, and their multi-group counterparts have not been extensively explored previously in the context of long-form text generation, we consider these empirical results to form a benchmark for this setting. To induce subgroups, we collect an additional set of attributes about the public figure (i.e., entity) for whom the LLM is producing a biography. We note that these groups are intersecting, meaning that each entity can belong to multiple subgroups. Our empirical results demonstrate that for both problems (1) and (2), multicalibration and multivalid conformal prediction techniques improve measures of uncertainty relative to standard (marginal) calibration and conformal prediction methods. This advantage holds regardless of whether evaluation is conducted within groups or across the entire dataset.
1.1 Related work
Factuality in long-form LLM outputs.
The task of evaluating factuality for long-form generation is challenging: not only do generated outputs consist of many parts that must be scored individually, but also scoring each part requires prohibitively costly manual annotation. To help make evaluation more tractable, Min et al. (2023a) introduce FActSCORE, which converts any generation into a set of atomic facts/claims that are then labeled true or false. Using this evaluation metric, Min et al. (2023a) test LLMs’ ability to generate biographies and find that their generations are pervaded with errors.
Attaching confidence scores to LLM outputs.
While the most natural method is to use a model’s output probabilities directly as a confidence (or uncertainty) score (Achiam et al., 2023), it has been shown that model probabilities are not-well calibrated (Guo et al., 2017). As a result, many works have recently proposed alternative methods for generating uncertainty scores that can then be used to refine or correct LLM outputs. Xiong et al. (2023), for example, explores whether a model can express uncertainty directly through prompting. Meanwhile, Wang et al. (2022) draw inspiration from self-consistency, in which a score is defined over how often an answer appears over multiple responses the same prompt. Our work is complementary to this line of work. Rather than proposing an entirely new uncertainty score function, we focus on how one can leverage existing scores to achieve uncertainty guarantees on LLM factuality.
Uncertainty quantification for LLMs.
Similar to Detommaso et al. (2024), we study multicalibration for LLM outputs. However, while Detommaso et al. (2024) calibrate for correctness in question-answering, we are the first to apply multicalibration to claims decomposed from long-form text generation. In addition, our work also closely relates to that of Mohri and Hashimoto (2024), who aim to provide high probability guarantees of factuality in long-form generation. In particular, Mohri and Hashimoto (2024) frame this problem as a nested conformal prediction problem, applying methods from works like Min et al. (2023a) to produce sets of claims outputted by an LLM that achieve some marginally valid coverage. In other words, their method produces some generation that on average, contains a correct output with any user-specified probability. Our work, however, extends this problem to multivalid conformal prediction: we produce generations that are not only correct on average but are also conditionally correct across subgroups.
2 Preliminaries
In this section, we describe calibration and conformal prediction, as well as their multi-group guarantee versions, multicalibration and multivalid conformal prediction.
2.1 Calibration
We begin by defining calibration in context of factuality in open-ended text generation. Suppose we are given some $(X,Y)\sim\mathcal{D}$ where $X\in\mathcal{X}$ denotes some claim outputted by an LLM, while $Y$ is an indicator in which $Y=1$ when the claim is correct (and $Y=0$ otherwise). Suppose there exists some uncertainty score function $f:\mathcal{X}\to[0,1]$ that measures confidence for the correctness of some input $X$ (with higher values denoting higher levels of confidence). Then a goal one may have when designing such a score function $f$ is to have that
$P_{\mathcal{D}}(Y=1\mid f(X)=p)=p,\forall x\in\mathcal{X}$ (1)
In other words, the probability that some LLM output is correct is given exactly by $f$.
Calibration, then, defines a simpler, more tractable condition, in which instead of ensuring guarantees across all possible values of $f$, we ensure a guarantee over coarser, level sets of our uncertainty score.
Definition 1.
(Calibration) A function $f$ is calibrated w.r.t $D$ if
$\Delta_{p}(f)=0,\forall p\in[0,1]$
where $\Delta_{p}(f)$ is the bias of $f$ for the p-th level set $S_{p}(f)=\{f(x)=p\}$:
$\Delta_{p}(f)=\mathbb{E}_{\mathcal{D}}[Y-f(X)\mid S_{p}(f)]$
Defining level sets is akin to dividing the output space of $f$ (i.e., $[0,1]$) into buckets. For example, one could round $f(X)$ to the nearest value in some predefined set of probabilities (e.g. $\{0,0.5,1.0\}$). One can view this definition of calibration as a desirable guarantee since it serves as a minimal condition for Equation 1—any $f$ that satisfies 1 must (at the very least) also be calibrated. We note that to evaluate calibration, we can consider the average squared calibration error (ASCE) of $f$.
$\textrm{ASCE}(f)=\mathbb{E}_{P}[\Delta^{2}_{P}(f)]$ (2)
The ASCE averages the squared bias across all level sets and is zero when $f$ is calibrated.
2.1.1 Multicalibration
While calibration provides an already important and useful guarantee, it can often be insufficient in many real-world scenarios. For example, in the context of generating information about people, one maybe desire that $f$ is calibrated not only across all people, but also within subpopulations defined by demographic attributes like sex or gender. Otherwise, it is possible that certain subgroups can still suffer from very high miscalibration, even when the score function is perfectly calibrated across $\mathcal{D}$. Ideally, one would hope to have guarantees while conditioning on as many subgroups in $\mathcal{X}$ as possible, both from the perspective of machine learning fairness as well as enhancing the likelihood of correctness in general.
Multicalibration (Hébert-Johnson et al., 2018) was developed to provide accurate guarantees across overlapping subgroups (i.e., a sample can belong to many groups). Let $g:\mathcal{X}\to\{0,1\}$ be a group function that evaluates to $1$ if $X$ belongs to some group. We study, then, the setting in which there exists of set of groups $\mathcal{G}$ that corresponds to our data domain $\mathcal{D}$. We note that while the set of groups can be disjoint, the problem of multicalibration then becomes trivial in this case because one can simply split a dataset into disjoint sets that can then each be calibrated individually. Consequently, prior work typically considers the more interesting case where many intersecting groups comprise $\mathcal{G}$
Given a group function $g$, we define group average squared calibration error (gASCE) as the following:
$\textrm{gASCE}(f,g)=\mathbb{E}_{P}[\Delta^{2}_{p,g}(f)\mid g(X)=1]$ (3)
where
$\Delta_{p,g}(f)=\mathbb{E}_{\mathcal{D}}[Y-f(X)\mid S_{p,g}(f)]$
for $S_{p,g}=\{f(X)=p,g(x)=1\}$. In other words, gASCE conditions on both level sets and group membership. Finally, we have:
Definition 2.
(Multicalibration) A function $f$ is $\alpha$-multicalibrated w.r.t $D$ and a set of groups $\mathcal{G}$ if and only if
$gASCE(f,g)<\frac{\alpha}{P_{\mathcal{D}}(g(X)=1)},\forall g\in\mathcal{G}$
2.2 Conformal Prediction
In conformal prediction, the general goal is to produce some confidence set $\mathcal{T}(X)$ for some example $X$ such that this set marginally covers the true label $Y$ with some target probability $1-\alpha$.
$P_{\mathcal{D}}(Y\in\mathcal{T}(X))=1-\alpha$ (4)
The second part of our work follows the problem statement outlined in Mohri and Hashimoto (2024). Unlike in calibration, where each claim contained in some long-form generation is treated individually, Mohri and Hashimoto (2024) instead define their problem in terms of pairs $(X,Y)$, where $X$ is some input prompt and $L(X)=Y\in\mathcal{Y}$ is the long-form generation outputted by a LLM $L$. Because $Y$ may or may not be supported by some reference ground truth $Y^{*}$,111e.g., in the case of FActScore (Min et al., 2023a), is $Y$ supported by Wikipedia. Mohri and Hashimoto (2024) define factuality in terms of entailment operations $Y^{*}\implies Y$. Furthermore, they rewrite this relation as $Y^{*}\in E(Y)=\{Y^{\prime}\in\mathcal{Y}:Y^{\prime}\implies Y\}$. This equivalent set notation, in other words, means that some reference ground truth $Y^{*}$ (e.g., Wikipedia article in Min et al. (2023a)) is contained in the set of possible texts $Y^{\prime}$ that support all claims made in the LLM output $Y$.
Given this notation, then, the goal is to find some uncertainty set $\mathcal{T}(L(X))$ s.t. $P_{\mathcal{D}}(Y\in\mathcal{T}(L(X)))=1-\alpha$. In the context of long-form text generation, this goal translates to taking as input the original LLM output $L(X)$ and producing a subset of claims $\mathcal{T}(L(X))$ such that with high probability, $1-\alpha$, all remaining claims are factually correct.
2.2.1 Multivalid Conformal Prediction
Similar to calibration, one may also desire group conditional coverage guarantees for intersecting groups. Known as multivalid conformal prediction (Jung et al., 2022), these guarantees are stronger than marginal conformal guarantees, holding also when conditioned on group membership. Using group functions $g$, as defined in Section 2.1.1, full multivalid coverage can be written as the following: Given some set of groups $\mathcal{G}$, we have that
$P_{\mathcal{D}}(Y\in\mathcal{T}(X)\mid g(X)=1)=1-\alpha$ (5)
for all group functions $g\in\mathcal{G}$.
3 Methods
Next, we introduce the methods (and their group-conditional variants) for applying calibration and conformal prediction to language model factuality. In particular, we organize these methods into two categories: (1) iterative "patching"-based algorithms and (2) linear regressor algorithms. As we mention previously, prior exploration of long-form text generation has been limited. While Mohri and Hashimoto (2024) evaluated one variant—split conformal (SC)—on a small set of entities, we are not aware of prior work that has considered other uncertainty quantification methods in this setting.
3.1 Iterative "patching" algorithms
The first category of algorithms can be characterized as "patching" algorithms. Given a base method for calibration or conformal prediction, one iterates through groups $g\in\mathcal{G}$ in which the method does poorly on. At each iteration, the patching algorithm corrects the bias (i.e., patches up the function) on just that subset of examples in the data domain (i.e., $g(x)=1$). Once some stopping condition is met, the final, "patched up" satisfies multi-group guarantees.
Calibration.
For calibration, we consider the standard method of Histogram Binning (HB) (Zadrozny and Elkan, 2001), presented in Algorithm 1. This method, takes some base scoring function $f$ and discretizes the output space to a set of $p$-th level sets $S_{p}(f)$, as defined in Section 2.1. Specifically, given some target grid of values $p\in{[\frac{1}{m}]}$, we round $f$ to the closest value in this grid
$f^{\prime}(x)=\operatorname*{argmin}_{p\in{[\frac{1}{m}]}}|f(x)-p|.$
Algorithm 1 then applies a constant correction [^2] In Algorithms 1 and 2, we assume that you are given the true data distribution and therefore can calculate $\Delta_{p,g}$. In practice, as is done for our experiments, $\Delta_{p,g}$ is estimated using a calibration set. for each level set $S_{p}(f)$ in the grid, based on the calibration error of the model $f^{\prime}$.
In Algorithm 2, we present the multi-group version of histogram binning, known as Iterative Grouped Histogram Binning (IGHB) (Hébert-Johnson et al., 2018). In this algorithm, we instead apply a constant correction conditioned on $S_{p,g}$ (i.e., both the level set and group membership). At each step $t$, IGHB identifies $S_{p,g}$ for which the calibration error (weighted by the group size) is highest and then corrects it for this level set and group. The algorithm then continues until some stopping condition is met. In other words, the method iteratively patches the output $f^{\prime}$ for various groups $g\in\mathcal{G}$.
Algorithm 1 Histogram Binning (HB)
1:  Input: scoring function $f^{\prime}$
2:  for $p\in[\frac{1}{m}]$ do
3:     Set
$\hat{f}(x)=\begin{cases}f^{\prime}(x)+\Delta_{p}(f^{\prime})&\text{if }x\in S_{p}(f^{\prime})\\ f^{\prime}(x)&\text{otherwise}\end{cases}$
4:  end for
5:  Output: $\hat{f}$
Algorithm 2 Iterative Grouped Histogram Binning (IGHB)
1:  Input: scoring function $f^{\prime}$, max iterations $T$
2:  Let $H_{t}(p,g)=P_{D}(S_{p,g}(f_{t}))\Delta^{2}_{p,g}(f_{t})$
3:  Initialize $f_{0}=f^{\prime}$
4:  for $t\in\{0,1,\ldots,T-1\}$ do
5:     Set
$(p_{t},g_{t})=\operatorname*{argmax}_{p\in\left[\frac{1}{m}\right],g\in\mathcal{G}}H_{t}(p,g)$
6:     Let $\Delta_{t}=\Delta_{p_{t},g_{t}}$ and $S_{t}=S_{p_{t},g_{t}}$
7:     Set
$h_{t+1}(x)=\begin{cases}f_{t}(x)+\Delta_{t}(f_{t})&\text{if }x\in S_{t}(f_{t})\\ f_{t}(x)&\text{otherwise}\end{cases}$
8:     Set $f_{t+1}=h_{t+1}^{\prime}$
9:     if $H_{t}(p_{t},g_{t})\geq H_{t-1}(p_{t-1},g_{t-1})$ then
10:        Set t = t-1
11:        break
12:     end if
13:  end for
14:  Output: $H_{t}$
Conformal prediction.
We first present the Split Conformal (SC) method (Shafer and Vovk, 2008; Gupta et al., 2022). In particular, we consider the standard approach where one constructs a set of nested sets and each output set contains some subset $\mathcal{F(X)}_{t}$ of claims generated by the LLM.
Following Mohri and Hashimoto (2024), we define these nested sets $\mathcal{T}$ as thresholds sets where each set $\mathcal{F}(L(X))$ contains the set of all individual claims $\{x\in L(X)\mid f(x)>t\}$ for some scoring function $f$ [^3] The same scoring function $f$ that one may wish to (multi)calibrate can also be used for split conformal. More formally, we have that $\mathcal{F}(L(X))_{t\in\mathcal{T}}$ satisfies the nested sequence property if for $t,t^{\prime}\in\mathcal{T},t\leq t$, we have that $\mathcal{F}_{t}(L(X))\subseteq\mathcal{F}_{t^{\prime}}(L(X))$.
To construct these threshold sets, we have the score
$r(X,Y)=\inf\{t\in\mathcal{T},Y\in\mathcal{F}_{t}(L(X))\}$
where $r$ defines the minimum safe threshold such that $Y\in\mathcal{F}_{t}(L(X))$ for all $t>r(X,Y)$. Practically speaking, given some set of uncertainty scores $f(x)$ for each claim $x\in L(X)$, $r(X,Y)$ defines the minimum value such that any set of claims $\mathcal{F}_{t}(L(X))=\{x\in L(X)\mid f(x)\geq t\}$ will be entirely true if and only if $t\geq r(X,Y)$.
Given some calibration set $\hat{D}$ of size $n$ and some target error rate $\alpha$ (or target coverage $1-\alpha$), split conformal simply outputs the set $\mathcal{F}_{q_{\alpha}}(L(X))$ for any $X$, where $q_{\alpha}$ is the $\frac{\lceil(n+1)(1-\alpha)\rceil}{n}$th-quantile of scores $\{r(X_{i},Y_{i})\}_{i=1}^{n}$ for $X_{i},Y_{i}\in\hat{D}$.
In Algorithm 3, we present the multivalid split conformal (MVSC) prediction technique that is closely related to methods originally proposed in Jung et al. (2022). Similar to IGHB, we start with some base threshold (i.e., the threshold $q_{\alpha}$ obtained from using split conformal). Then at each iteration $t$, we find the group $g_{t}$ that has the worst squared coverage error $\Delta_{t,g}$, weighted by the size of the group $P(g_{t}(X)=1)$. Then, we simply "patch" the thresholds for examples $\{(X,Y)\mid g_{t}(X)=1\}$, again using the $\frac{\lceil(n+1)(1-\alpha)\rceil}{n}$th-quantile of scores for $(X,Y)$ beloing to group $g_{t}$. Like in IGHB, we continue patching the set of thresholds until some stopping criterion is met.
Algorithm 3 Multivalid Split Conformal (MVSC)
1:  Input: calibration set $\hat{D}$, LLM $L$, fact-level scoring function $f$, target error rate $\alpha$, split conformal threshold $q_{\alpha}$, max iterations $T$
2:  Let $\mathcal{F}_{h_{t}}(L(X))=\{x\in L(X)\mid f(x)\geq h_{t}(X)\}$
3:  Let $\Delta_{t,g}=P_{\mathcal{D}}(Y\in\mathcal{F}_{h_{t}}(L(X))\mid g(X)=1)$
4:  Let $H_{t}(g)=P_{\mathcal{D}}(g(X)=1)[(1-\alpha)-\Delta_{t,g}]^{2}$
5:  Initialize $h_{0}(X)=q_{\alpha}$
6:  for $t\in\{0,1,\ldots,T-1\}$ do
7:     Set
$g_{t}=\operatorname*{argmax}_{g\in\mathcal{G}}H_{t}(g)$
8:     Let $\hat{D}_{t}=\{(X,Y)\in\hat{D}\mid g_{t}(X)=1\}$
9:     Set $q_{t}$ to be the $\frac{\lceil(n+1)(1-\alpha)\rceil}{n}$th-quantile of scores $\{r(X_{i},Y_{i})\}$ for $X_{i},Y_{i}\in\hat{D}_{t}$
10:     Set
$h_{t+1}(X)=\begin{cases}q_{t}&\text{if }g_{t}(X)=1\\ f_{t}(X)&\text{otherwise}\end{cases}$
11:     if $H_{t}(g_{t})\geq H_{t-1}(g_{t-1})$ then
12:        Set t = t-1
13:        break
14:     end if
15:  end for
16:  Output: $h_{t}$
3.2 Linear regressor algorithms
Next, we consider a set of algorithms that instead solve some optimziation problem for the purpose of calibration and conformal prediction. In these cases, one can naturally make them multi-group/valid by including group-membership (i.e., $g(X)=1$ for all $g\in\mathcal{G}$) in the optimization problem itself. Formally, we describe these linear regression based methods in Algorithms 4 and 5. Presented in this way, the methods for calibration vs. conformal prediction is reduced to a choice of loss function $L$. Again, we assume one has access to some calibration set for which one solves the optimization problem on.
Calibration.
In the case of calibration, one can choose $L$ to be binary cross-entropy loss. In doing so, Algorithm 4 then describes Platt Scaling (PS) (PLATT, 1999), which can be described as fitting a logistic regression model to some set of model outputs to obtain calibrated probability scores. Algorithm 5 describes the multicalibrated version of Platt Scaling (Gopalan et al., 2022). Going forward, we refer to this method as Group Conditional Unbiased Logistic Regression (GCULR).
Conformal prediction.
For conformal prediction, we instead choose $L$ to be pinball loss. We refer to the non-group version of this method (Algorithm 4) as Conformalized Quantile Regression (CQR) (Romano et al., 2019), in which given some target coverage $1-\alpha$, we fit a linear quantile regression model that minimizes pinball loss.
In our conformal prediction setting, as described in Section 2.2, $X$ is an entire biography, or set of independent claims. Thus, to adapt quantile regression to long-form generation, we propose setting $f(X)$ to be a vector of uncertainty scores for each claim $x\in X$ [^4] Since each biography generated by the LLM may have a different number of claims, we apply interpolation to (un)squeeze the set of scores to a list of $K=25$ scores. Like in split conformal, the target is then the minimum threshold $r(X,Y)$ for which all claims above it are correct. In the multivalid case, we then add group features $g(X)$ to the optimization problem. A version of Algorithm 5 was first presented by Jung et al. (2022), and going forward, we will refer to this method as Group Conditional Conformalized Quantile Regression (GCCQR). Again, we set $f(X)$ be some vector of (interpolated) scores.
Algorithm 4 Linear Regressor
1:  Input: scoring function $f$, loss function $L$
2:  Set
$\hat{\lambda}=\operatorname*{argmin}_{\lambda}\mathbb{E}_{(X,Y)\sim\mathcal{D}}\left[L\left(f(X;\lambda),Y\right)\right]$
s.t. $f(X;\lambda)=\lambda_{0}+\lambda_{1}f(X)$
3:  Output:$f(X;\hat{\lambda})$
Algorithm 5 Group-conditional Linear Regressor
1:  Input: scoring function $f$, loss function $L$, set of groups $\mathcal{G}$
2:  Set
$\hat{\lambda}=\operatorname*{argmin}_{\lambda}\mathbb{E}_{(X,Y)\sim\mathcal{D}}\left[L\left(f(X;\lambda),Y\right)\right]$
s.t. $f(X;\lambda)=\lambda_{0}+\lambda_{1}f(X)+\sum_{\lambda_{g}\in\mathcal{G}}\lambda_{g}g(x)$
3:  Output:$f(X;\hat{\lambda})$
4 Empirical evaluation
Model Metric Uncal. HB IGHB PS GCULR
Llama 2 7B Chat marginal 0.32291 0.00875 0.00038 0.00022 0.00015*
group max 0.42343 0.07711 0.01481 0.05791 0.00628*
group mean 0.33352 0.01597 0.00289 0.00636 0.00111*
Mistral 7B Instruct marginal 0.30706 0.01163 0.00026 0.00029 0.00013*
group max 0.43659 0.08372 0.01660 0.05729 0.00487*
group mean 0.32067 0.01988 0.00269 0.00726 0.00097*
Table 1: We generate biographies for entities from Bio-NQ. We evaluate using ASCE, max gASCE, and average gASCE ($\downarrow$ better), comparing each calibration method (HB, PS) against its multicalibration counterpart (IGHB, GCULR). We bold the better-performing method for each pairing. * denotes the best-performing method across all methods evaluated.
Model Metric Uncal. HB IGHB PS GCULR
Llama 2 7B Chat marginal 0.463 0.218 0.190 0.196 0.181*
group max 0.574 0.293 0.260 0.267 0.247*
group mean 0.470 0.208 0.179 0.186 0.171*
Mistral 7B Instruct marginal 0.455 0.227 0.194 0.201 0.187*
group max 0.609 0.302 0.258 0.271 0.245*
group mean 0.466 0.216 0.180 0.190 0.174*
Table 2: We generate biographies for entities from Bio-NQ. We evaluate using Brier score ($\downarrow$ better) marginally across the entire dataset, as well as within each subgroup (in terms of max and mean overall all groups). We compare each calibration method (HB, PS) against its multicalibration counterpart (IGHB, GCULR). We bold the better-performing method for each pairing. * denotes the best-performing method across all methods evaluated.
We focus our empirical evaluation on the problem of biography generation. Following Min et al. (2023a), we use an LLM to automate the process of decomposing biographies into individual claims and evaluating for factuality. Details can be found in Appendix A.
4.1 Dataset
We evaluate on a large set of biographies by extracting 8,541 entities from Natural Questions dataset (Kwiatkowski et al., 2019), which consists of real queries issued to the Google search engine. We denote this dataset as Bio-NQ. Our motivation for choosing Natural Questions is that these extracted human entities should serve as a representative sample of public figures that users may prompt an LLM to generate information for. To form our dataset, for each question, we select all entities in either the question’s short answer or accompanying Wikipedia article. We then attempt to match them to their corresponding Wikidata entry. If a match exists and its Wikidata page’s property, if instance of, is equal to the value, human, we add the entity to our dataset.
4.2 Collecting group features
To obtain groups for each person found in our dataset, we extract properties by scraping Wikidata for each entity and identifying ones that are commonly shared among entities in Bio-NQ. For our experiments, we use the following group attributes:
  • # Wikidata properties: For each entity, we count the number of Wikidata properties and discretize them into the following buckets: [0,25),[25,50),[50100),[100,)[0,25),[25,50),[50-100),[100$\infty$)[ 0 , 25 ) , [ 25 , 50 ) , [ 50 - 100 ) , [ 100 , ∞ ). This group serves as proxy for the amount of information available online for some given entity.
  • nationality: Following Min et al. (2023a), who use nationality derived from Wikidata to sample their dataset of human entities, we take the property country of citizenship (or place of birth when not available) and categorize the corresponding value into the following categories defined by Min et al. (2023a): Asia/Pacific, Europe/Middle East, North America, Latin/South America/Africa.
  • sex or gender: We take directly the value for the Wikidata property, sex or gender.
  • plays professional sports: We check whether the Wikidata entry has the property, sport.
  • has IMDb entry: We check whether the Wikidata entry has the property, IMDb ID, to use as a proxy for whether a person has been involved in films or television series.
To form groups $\mathcal{G}$ from these attributes, we take all $1$ and $2$-way combinations [^5] To reduce computation, we train the quantile regression models in CQR and GCCQR using only single attribute groups as features in the optimization problem, excluding 2-way combinations. of attributes and the values they take on. In total, we have $|\mathcal{G}|=77$ subgroups. [^6] To prevent extremely uncommon groups that may exist in the Wikidata database from biasing our results, we exclude groups of size $<5\%$ of the total test set size.
4.3 Generating confidence scores
As mentioned previously, the algorithms described in Section 3 require a base scoring function. For our experiments, we use a frequency-based scoring function derived from concept of self-consistency (Wang et al., 2022). Specifically, to score each claim found in a generated biography, we prompt the LLM to output a biography $M$ additional times. We use the proportion of times the claim is contained in the additional reference generations as the uncertainty score. To automate this process, we use a procedure similar to frequency scoring algorithm proposed by Wang et al. (2024) in which (1) a set of $K$ most relevant claims from a reference generation is retrieved using a vanilla BM25 algorithm (to reduce computational costs). Then (2) an LLM is tasked to evaluate whether the target claim is supported by the set of $K$ reference claims. In our work, we replace LLM prompting in step (2) with AlignScore-Large (Zha et al., 2023), which runs significantly faster and is reported by Zha et al. (2023) to compare favorably to LLM-based alignment methods.
5 Results
To assess the efficacy of these methods for quantifying the uncertainty of LLM factuality, we present empirical results for the task of biography generation, using outputs from Llama 2 7B Chat (Touvron et al., 2023) and Mistral 7B Instruct v0.2 (Jiang et al., 2023). We randomly split the entities into 80-20 calibration-test splits. The results we present are averaged over 10 distinct randomly generated splits. Table 5 reports the number of entities for which each model generates biographies and the number of claims contained in each generated biography.
Calibration.
To evaluate fact-level uncertainty, we consider both the ASCE and the Brier score, which is the mean squred error between the uncertainty score function $f(X)$ and the true label $Y$. While not a direct measure of (multi)calibration like the ASCE, the Brier score is still useful in certain settings for quantifying the efficacy of the algorithms we consider. As discussed previously, we generate biographies for entities from Bio-NQ and apply the techniques discussed in Section 3.
In Table 1, we report the ASCE, max gASCE, and mean gASCE, comparing each calibration method (HB, PS) against its multicalibration counterpart (IGHB, GCULR). The multicalibration variants of both the patching and linear regression techniques significantly outperform HB and PS in terms of max and mean gASCE. More surprising, however, is that even when considering just the ASCE across the entire dataset, incorporating group features improves performance as well, especially when comparing IGHB to HB. These results suggest that even if one does not specifically require parity for specific subgroups, collecting additional group features and applying multicalibration (as opposed to vanilla calibration) can still be extremely beneficial for generating better-calibrated uncertainty scores. In Table 2, we report Brier score, also both marginally and across groups. Again, we observe that IGHB and GCULR outperform HB and PS respectively across all metrics.
Conformal prediction.
For the problem of uncertainty at the biography level, we apply the vanilla conformal prediction methods SC and CQR and their multivalid counterparts, MVSC and GCCQR. We choose target coverages of between $0.5$ to $0.9$. [^7] Although we evaluate on a wide set of target coverages $1-\alpha$, conformal prediction makes more sense only for higher target coverages (e.g., $0.8$ or higher), since lower coverage guarantees can often be too weak to be useful in practice. We initially observe that for both Llama 2 7B Chat and Mistral 7B Instruct, all methods are able to achieve close to perfect coverage on Bio-NQ (Figure 2, left panel). However, when evaluating coverage across individual subgroups, we find that all methods have some level of error. In Figure 1, we compare the mean absolute coverage across all subgroups for each target coverage. Like for calibrating individual facts, we again find that multivalid variants of both the patching (MVSC) and linear regression (GCCQR) techniques outperform standard conformal techniques (SC, CQR).
(a) LLama 2 7B Chat
(b) Mistral 7B Instruct
Figure 1: For each target coverage $1-\alpha$, we run conformal prediction methods (SC, CQR) and their multigroup counterparts (MVSC, GCCQR). On the top row of plots (a), we evaluate on biographies generated by Llama 2 7B Chat, and on the bottom row (b), we evaluate on biographies generated by Mistral 7B Instruct. We calculate the average coverage error across all subgroups and plot them side by side for each pair of methods.
6 Conclusion
In this paper, we conduct an extensive study on uncertainty quantification in long-form LLM generation. We focus on two forms of uncertainty quantification for factuality in long-form generation—claim-level (calibration) and biography-level (conformal prediction)—and present a variety of methods for these settings. Introducing two categories of algorithms (iterative patching and linear regression), we demonstrate that by accounting for additional groups, multicalibration and multivalid conformal prediction methods outperform their marginal-guarantee counterparts across group-wise performance and, in the case of calibration, overall marginal performance. We consider these empirical results to establish a benchmark for this setting and hope our findings will motivate future work in quantifying uncertainty for LLMs.
7 Limitations and potential risks
Past works on uncertainty quantification for NLP have made great strides in providing confidence scores for text generation, and our work furthers this line of research. However, we note that there exists certain limitations in our empirical findings. First, while Min et al. (2023a) demonstrate the effectiveness of using an LLM to both decompose generations into atomic facts and assess the factuality of these claims, this automated process serves only as a proxy for the gold standard of human annotation. Moreover, like in Min et al. (2023a), our assessment of factuality is limited to whether claims are supported by Wikipedia, which again serves only as a proxy for factuality.
In our experiments, we chose groups based on attributes that are general and easily obtainable for any public figure (i.e., Wikidata properties). This set of groups is non-exhaustive, and there may be many other sets of groups one can consider that may be relevant for various problem domains.
Finally, while we follow prior work and focus on the problem of generating biographies, our uncertainty techniques extend broadly to any long-form generation task where information contained in the output can be broken down into independent claims. Furthermore, the automated evaluation pipeline proposed by Min et al. (2023a) is still valid, as long as the entities one prompts the LLM for a description of can be found on Wikipedia (i.e., animals, places, historical events, etc.). Consequently, in future work, we hope to validate our findings for other long-form generation tasks.
Regarding potential risks, as mentioned previously, large language models can have a tendency to hallucinate. Moreover, deep learning models in general may exhibit or perpetuate bias. In an attempt to better quantify uncertainty of LLMs, our work aims to help alleviate the potential risks of models producing false information. Moreover, our emphasis on multi-group guarantees aligns with certain definitions of fairness. However, our work may not entirely mitigate such issues, and other potential malicious or unintended harmful effects could still persist in the outputs of the calibration and conformal methods we evaluate.
Appendix A Details on biography generation and factuality evaluation
While the ground truth score must be human-annotated, Min et al. (2023a) show that FActSCORE can be approximated by an automated process that leverages an LLM (i.e., ChatGPT and LLaMa-7B) and natural language retrieval. Following Min et al. (2023a), we also use an LLM to automate the annotation process. For some input person, [ENTITY], we prompt a large language model with the following:
[INST] Question: Tell me a bio of [ENTITY]. [/INST]
We then decompose each long-form generation into a set of atomic facts, which are then checked against some set of Wikipedia articles about the [ENTITY] to evaluate overall performance of language model in terms of factuality. Min et al. (2023a) demonstrate that while the evaluation process should ideally be conducted by human annotators, using large language models (i.e., ChatGPT and LLama 1 7B) to both decompose long-form generations and check against Wikipedia articles serves as a very good proxy for human annotation.
Following this general framework for automated evaluation, we use Llama 2 7B Chat to decompose each generation [GEN_BIO] with the following prompt:
[INST] <<SYS>> Break down the following input into a set of small, independent claims. You must not add additional information. Output the claims as a numbered list separated by a new line. The subject of each line should be [ENTITY]. <</SYS>> Input: [GEN_BIO] [/INST]
For checking each atomic fact against Wikpedia, we directly use the code released by Min et al. (2023a), which first conducts passage retrieval via Generalizable T5-based Retrievers (Liu et al., 2023) to find relevant articles from a dump of Wikipedia (dated 2023-04-01) and then prompts an LLM (i.e., ChatGPT or Llama 1 7B) to predict whether each fact is supported by the retrieved passages. For our evaluation, we again use Llama 2 7B Chat. Finally, these predictions are ensembled with predictions using likelihood estimates derived from a nonparametric masked language model (Min et al., 2023b).
We note that for prompting the LLMs described above, we use Hugging Face’s transformer’s library and generate responses with temperature set to $1.0$.
Appendix B Additional results
B.1 Additional results on Bio-NQ
In Figure 2, we provide additional information about the prediction sets outputted by our various conformal methods on Bio-NQ. On the left panel, we plot the empirical coverage achieved against the target coverage. Figure 2 demonstrates that all methods achieve the target (marginal) coverage. On the middle panel, we plot the fraction of biographies retained for each method against the target coverage level, while on the right panel, we plot the number of facts per biography retained. Generally speaking, all methods retain about the same number of facts per biography. We also observe that to achieve the same target coveragre, SC and MVSC generally retain fewer biographies (i.e., more abstentions) when compared to CQR and GCCQR. However, when comparing each conformal method (SC and CQR) to their multivalid counterparts (MVSC and GCCQR), we again observe that there are very little differences between them.
(a) LLama 2 7B Chat
(b) Mistral 7B Instruct
Figure 2: We report additional metrics for conformal predictions techniques when evaluated on biographies generated for Bio-NQ: On the top row, we present these metrics for outputs from LLama 2 7B Chat, and on the bottom, Mistral 7B Instruct. On the left panel, we plot the empirical coverage achieved against the target coverage. On the middle panel, we plot the fraction of biographies retained for each method against the target coverage level. Finally, on the right panel, we plot the number of facts per biography retained, again against the target coverage level.
B.2 Evaluating on entities used in Min et. al (2023a)
In addition to evaluating on our dataset, Bio-NQ, we construct an additional dataset using the $683$ entities used by Min et al. (2023a) for their empirical evaluation. We denote this dataset as Bio-FActScore.
Calibration.
In Table 3, we observe similar results to that on Bio-NQ—namely, multicalibrated counterparts (IGHB and GCULR) perform better than their base variant (HB and PS). However, we note that for Mistral 7B Instruct, PS performs the best when looking at marginal ASCE. We hypothesize that the smaller gap in ASCE between PS and GCULR may be due to the smaller training size of Bio-FActScore (25,283 claims), which is roughly 10x smaller than that of Bio-NQ (297,714 claims). Lastly, with respect to Brier score, multicalibration still dominates across all metrics (Table 4).
Model Metric Uncal. HB IGHB PS GCULR
Llama 2 7B Chat marginal 0.26830 0.00951 0.00229 0.00164 0.00125*
group max 0.48594 0.07208 0.04088 0.05017 0.03519*
group mean 0.29983 0.02848 0.01108 0.01659 0.00858*
Mistral 7B Instruct marginal 0.25496 0.01032 0.00268 0.00093* 0.00146
group max 0.54701 0.08436 0.04585* 0.07043 0.04931
group mean 0.29435 0.03226 0.01143 0.01848 0.00911*
Table 3: We generate biographies for entities from Bio-FActScore. We evaluate using ASCE, max gASCE, and average gASCE ($\downarrow$ better), comparing each calibration method (HB, PS) against its multicalibration counterpart (IGHB, GCULR). We bold the better performing method for each pairing. * denotes the best performing method across all methods evaluated.
Model Metric Uncal. HB IGHB PS GCULR
Llama 2 7B Chat marginal 0.475 0.169 0.148 0.152 0.143*
group max 0.535 0.323 0.247 0.285 0.235*
group mean 0.479 0.169 0.148 0.152 0.142*
Mistral 7B Instruct marginal 0.471 0.186 0.159 0.164 0.152*
group max 0.554 0.333 0.250 0.285 0.235*
group mean 0.477 0.186 0.158 0.164 0.152*
Table 4: We generate biographies for entities from Bio-FActScore. We evaluate using Brier score ($\downarrow$ better) marginally across the entire dataset, as well as within each subgroup (in terms of max and mean overall all groups). We compare each calibration method (HB, PS) against its multicalibration counterpart (IGHB, GCULR). We bold the better performing method for each pairing. * denotes the best performing method across all methods evaluated.
Conformal Prediction.
For Bio-FActScore, we observe that multivalid conformal methods do not improve performance across subgroups. In Figure 3, we observe very little difference in mean coverage error across groups. We hypothesize, however, that this negative result again is due to the smaller dataset size. In this case, our number of examples is the number of biographies in the dataset (683), giving us a calibration set size of 546 and test set size of 137. Further dividing the calibration and test sets into subgroups, it is possible there could simply not be enough examples per group for the distribution on the calibration set to generalize to the test set. Comparing the left panels of Figures 4 to 2, we also find that even when looking at marginal coverage, all methods perform worse (the lines deviate more from $y=x$), likely due again to the small calibration and test size.
(a) LLama 2 7B Chat
(b) Mistral 7B Instruct
Figure 3: For each target coverage $1-\alpha$, we run conformal prediction methods (SC, CQR) and their multigroup counterparts (MVSC, GCCQR). On the top row of plots (a), we run the methods on biographies generated by Llama 2 7B Chat, and on the bottom row (b), we run on biographies generated by Mistral 7B Instruct. We calculate the average coverage error across all subgroups and plot them side by side for each pair of methods.
(a) LLama 2 7B Chat
(b) Mistral 7B Instruct
Figure 4: We report additional metrics for conformal predictions techniques when evaluated on biographies generated for Bio-FActScore: On the top row, we present these metrics for outputs from LLama 2 7B Chat, and on the bottom, Mistral 7B Instruct. On the left panel, we plot the empirical coverage achieved against the target coverage. On the middle panel, we plot the fraction of biographies retained for each method against the target coverage level. Finally, on the right panel, we plot the number of facts per biography retained, again against the target coverage level.
Appendix C Additional experimental details
Code for reproducing all results will accompany the final version of this work.
Datasets.
Table 5 reports additional information about datasets Bio-NQ and Bio-FActScore, including the number of entities and claims per biography outputted by each model.
dataset model # entities total # claims avg. # claims per bio.
Bio-FAcTScore Llama 2 7B Chat 683 17,605 25.78
Mistral 7B Instruct 683 25,283 37.02
Bio-NQ Llama 2 7B Chat 8,541 206,620 24.19
Mistral 7B Instruct 8,541 297,714 34.86
Table 5: We provide statistics about our two datasets and how many claims are generated by each LLM.
Hyperparameters.
For our patching algorithms IGHB and MVSC, we set the max iterations $T=100$. For training (multi)calibration, our logistic regression models are trained using default hyperparameters given my Sci-kit learn. For training CQR and GCCQR, we run 5-fold cross validation for each target coverage $1-\alpha$ to optimize the $\ell_{1}$-penalty term $C\in\{10^{-6},10^{-5},10^{-4},10^{-3},10^{-2},10^{-1}\}$
For AlignScore, we set $M=4$ and $K=5$. We found that AlignScore generally returns values close to 00 or $1$, giving us uncertainty scores around the $5$ values $\{0,\frac{1}{4},\frac{1}{2},\frac{3}{4},1\}$. As a result, we evaluate all methods using $p=5$ level sets. We note that instead of rounding to $\{0,\frac{1}{4},\frac{1}{2},\frac{3}{4},1\}$, we use a $k$-means clustering algorithm, where $k=5$, on the calibration set to determine what values to round to for IGHB (after each iteration), PS, and GCULR. Note that this method of rounding is not applicable to HB.
GPU requirements.
We use a NVIDIA A100 80GB GPU for all experiments. For obtaining results on all entities across Bio-NQ and Bio-FActScore, our experiments, per LLM require approximately the following:
  • Generating biographies (+ 4 additional generations for getting frequency scores): 15 hours (x5)
  • Splitting atomic facts (+ 4 additional generations for getting frequency scores): 30 hours (x5)
  • Checking facts against Wikipedia: 75 hours (x1)
  • Calculating frequency scores via AlignScore: 10 hours (x1)
Licenses.
Wikidata and Wikipedia are licensed under the Creative Commons CC0 License. Llama 2 7B is licensed under Meta’s Llama 2 license. Mistral 7B Chat and Hugging Face’s transformers library are licensed under Apache 2.0 license. We also make use of code released by Min et al. (2023a) under the MIT license.